# A Monte Carlo Tree Search approach to QAOA: finding a needle in the haystack

### v2

@@ Abstract
The search for quantum algorithms to tackle classical combinatorial optimization problems has long been one of the most attractive yet challenging research topics in quantum computing. In this context, variational quantum algorithms (VQA) are a promising family of hybrid quantum-classical methods tailored to cope with the limited capability of near-term quantum hardware. However, their effectiveness is hampered by the complexity of the classical parameter optimization which is prone to getting stuck either in local minima or in flat regions of the cost-function landscape. The clever design of efficient optimization methods is therefore of fundamental importance for fully leveraging the potential of VQAs. In this work, we approach parameter optimization as a sequential decision-making problem and tackle it with an adaptation of Monte Carlo Tree Search (MCTS), a common artificial intelligence technique designed for efficiently exploring complex decision graphs. We show that leveraging regular parameter patterns deeply affects the decision-tree structure and allows for a flexible and noise-resilient optimization strategy suitable for near-term quantum devices. Our results shed further light on the interplay between artificial intelligence and quantum information and provide a valuable addition to the toolkit of variational quantum circuits.
@@ I Introduction
Overcoming the limitations of classical computing for complex classical problems has been one of the main drives behind the development of quantum information theory and computation since the first pioneering works on quantum algorithms. Indeed, some celebrated examples, such as Grover’s [1] and Shor’s [2] algorithms, display significant speedup over classical methods, motivating further research in the quest for quantum advantage. However, the realization of fault-tolerant quantum computers, a prerequisite for fully exploiting this potential, remains a formidable challenge. In the interim, leveraging available technology necessitates exploring alternative strategies, such as variational quantum algorithms [3, 4, 5, 6] (VQA), to harness the computational power of current quantum devices.
@@
Despite the flexibility and popularity of VQAs, one of their main weaknesses appears in the classical optimization loop [33], which may suffer from hardware and measurement noise or the presence of flat regions in the cost function, known as barren plateaus [34]. Indeed, both effects either slow down the exploration of parameter space or prevent it from escaping local minima, in particular when the classical optimization algorithm relies on the gradient of the cost function. Hence, gradient-free approaches represent an appealing alternative for the practical implementation of VQAs. Among these, reinforcement-learning [35] inspired approaches emerged as intriguing choices, both for the parameter optimization [27, 28, 36] and circuit design [30, 37].
@@
[…]
@@
The rest of the paper is organized as follows. In section II we recap the general structure of MCTS algorithms (II.1), describe the MCTS approach for QAOA (II.2) and explain the SSR variant and its iterative version (II.3). In sections III and IV we present our results on instances of 3-SAT and unweighted MaxCut on 3-regular graphs, respectively. We draw our conclusion and present future outlooks in Sec. V. Further results and analysis are presented in the appendices: in App. A we briefly discuss a combination of MCTS and gradient descent to construct a global optimization algorithm, in App. B we look at an example MCTS decision tree for QAOA, in App. C we test the performance of various iterative SSR-MCTS variants on noisy cost landscapes, and in App. D we study the resource cost of our algorithm.
@@ II.1.1 General concepts
MCTS explores a given tree domain, in which one or multiple agents interact, via random samples that are guided by some specified selection and playout policies, in order to find a path down the decision tree that maximizes a reward associated with the task at hand [38, 40]. Starting from the tree’s root node, MCTS incrementally builds an asymmetric replica of the parts of the tree (branches) it deems most promising. The algorithm repeatedly runs a four-stage cycle, learning about the tree by growing the part stored in memory. At the same time, it keeps track and continuously updates the cumulative score $w_{i}$ of each stored node $i$ and the number of times it has been visited by the algorithm $n_{i}$. The cycles stop when the allotted computational budget is exceeded, at which point the algorithm makes a decision and suggests a strategy to follow. This usually corresponds either to the child node with the maximum average score $w_{i}/n_{i}$, or the most visited one by the algorithm during the cycles. The larger the number of cycles, the more complete the reconstruction of the tree will be, leading to more accurate statistics of the nodes and improving the quality of the final choice. If the number of cycles is unbounded, MCTS is eventually equivalent to a brute force search and the policy converges to the optimal strategy [41].
II.1.2 The MCTS cycle
@@
    In the selection stage, starting from the root, the algorithm travels the part of the tree stored in memory by selecting for each node its child that maximizes a specified selection policy. The most common choice is the UCB1 policy [38, 48], which selects node $a^{*}$ among the children of node $i$, $A(i)$, such that
    $a^{*}=\operatorname*{arg\,max}_{a\in A(i)}\left\{\frac{w_{a}}{n_{a}}+C\sqrt{\frac{2\ln n_{i}}{n_{a}}}\right\},$ (2)
@@ II.2 MCTS for quantum optimization
The use of MCTS for QAOA parameter optimization requires formulating the task as a sequential decision problem. The most straightforward way to achieve this is to build a decision tree of depth $2P$ and branching factor $b$, where at each depth $i\in[1,2P]$ the algorithm must choose a value for the parameter $\theta_{i}$, where $\bm{\theta}=\{\gamma_{1},\beta_{1},...,\gamma_{P},\beta_{P}\}$, out of $b$ discrete options, $\theta_{i}\in\{\theta_{i}^{(1)},...,\theta_{i}^{(b)}\}$, within a range which depends on the spectrum and the symmetries of the Hamiltonians generating the parameterized unitary operators. As we consider Hamiltonians with an integer spectrum, the variational parameters belong to the compact interval $[0,2\pi)$. Moreover, since the cost function is real, it satisfies $F_{P}(-\bm{\gamma},-\bm{\beta})=F_{P}(\bm{\gamma},\bm{\beta})$, meaning that, if the different $\theta_{i}^{(k)}$ are uniformly distributed, for every path $\bm{\theta}$ there is an equivalent one where $\theta_{i}^{(k)}\rightarrow 2\pi-\theta_{i}^{(k)}$. Hence, we can restrict the domain of the first parameter to $[0,\pi]$, effectively halving the size of the tree to a total of $\left\lceil\frac{1}{2}b^{2P}\right\rceil$ leaf nodes, without changing the discretization step.
This approach transforms a continuous optimization problem into a discrete one, however, we do not expect it to affect QAOA performances significantly. Indeed, the algorithm’s efficiency is limited by the complexity of the energy landscape and the structure of the decision tree, not by the impossibility of exactly pinpointing the target minimum because it does not fall on the discrete grid. Yet, pathological cases might exist where extremely thin minima [20, 53] and wide flat regions require more clever discretization schemes than the regular ones.
@@
As we rely on the variational principle for the optimization, the reward collected at each leaf node must be related to the energy associated with the corresponding choice of parameters $F_{P}(\bm{\gamma},\bm{\beta})$. In this work, we chose a nonlinear mapping $r(\bm{\theta})=e^{-\nu F_{P}(\bm{\theta})}$ for the rewards, with $\nu$ chosen such that the reward starts rising when energies at or below the typical ranges are considered, and the algorithm can properly gauge the goodness of the parameters in the context of the problem at hand  [^1] To ensure the theoretical bounds on the regret of the playing agent [94], the UCB1 selection policy with $C=\sqrt{2}$ requires the rewards to lie in [0,1], although this is not strictly necessary for MCTS to perform well, and should be studied on a case-by-case basis [50]..
In MCTS language, optimizing the QAOA parameters is a single-player game, meaning that no opposing agent is trying to reach a different goal. This has some consequences for the optimal strategy choice: in a two-player game, a meaningful strategy is based on the assumption that the opponent is trying to win as well, and high-scoring paths based on counterproductive choices of the other agent are usually not worth exploring [41]. This is not true in a single-player game, where the goal is maximizing the score considering all eligible strategies. In view of this, we implement some changes to MCTS proposed in Ref. [55]. We first modify the criterion with which MCTS decides the final move to play, by choosing the child that leads to the path of the random simulation that obtained the highest reward. Second, we expand the highest-scoring path of each turn all the way to the leaf node in between turns, thus ensuring that MCTS memorizes such good paths in their entirety. The search still follows the UCB1 policy in (2) for the selection stage, which can under or overestimate the potential of exploration of certain regions of the search domain. But with these changes, we have the certainty that the final parameters MCTS chooses will be at least as good as those in the best memorized sequence. In the following, we will denote this modified variant as MCTS-SP (single-player).
In the following sections, however, we will still consider the standard version of MCTS. First, it is interesting to see how it compares to the single-player variant and, second, we expect the single-player modifications to no longer be beneficial in some particularly relevant cases, such as when considering a noisy cost landscape (see App. C). In this noisy situation, the task loses its deterministic nature, and the memorized parameter sequences will likely contain inaccurate expected rewards, making the standard final choice criteria more appropriate, and rendering path memorization irrelevant.
@@ II.3 Iterative Search-Space Restricted MCTS
MCTS has been proven to work well for optimizing annealing schedules [36, 42], suggesting that similar performances are expected when applied to QAOA, at least for shallow depths. In our study, we found that standard MCTS is unsuited for this purpose as soon as $P>1$. The modified single-player version obtains mildly improved results but quickly fails to learn good strategies too, even in shallow circuits, as we will discuss later on in Sec. III. Therefore, MCTS needs to be adapted further to perform well in QAOA parameter optimization, leveraging on the structure of its energy landscape. In this section, we describe how to modify the standard MCTS algorithm to successfully tackle QAOA.
@@
The core idea is not to modify the MCTS algorithm per se, but to progressively modify the space of eligible strategies. We call this method iterative search-space restriction (SSR). It consists of constraining the action space of the decision tree nodes as the circuit depth $P$ increases, according to the knowledge of previous solutions for $P-1$. Indeed, QAOA optima often display regular structures and patterns in the optimal parameters [9, 19, 28, 46], look ahead to Fig. 5 for an example. This property stems from the presence of optimal discretized quantum annealing schedules, which are typically robust with respect to small parameter changes in the Hamiltonian, system size, and number of time steps. These regular patterns appear as a monotonic increase (decrease) of the $\gamma_{i}$ ($\beta_{i}$) with the discrete-time index $i$, approaching a continuous function as $P\to\infty$. Hence, when increasing the number of layers, the new optimal parameters typically resemble a finer interpolation of the optimal schedule for the previous circuit depth.
@@
This method can be employed to look for the optimal QAOA angles one layer deeper when we know the optima for the previous depth. Additionally, the iterative application of SSR starting from $P=1$, allows for a systematic restriction of the action space by gradually increasing the number of parameters, without any prior knowledge of the optimal strategy. This strategy consists in making use of the parameters suggested by SSR-MCTS at depth $P$ to restrict the $P+1$ search. The different $\delta_{P}$ for each step can be adjusted to relax the restriction and help the algorithm correct its potential inaccuracies in previous circuit depths. In this iterative case, the boundaries of the first and last parameters $k=0$ and $k=P+1$ do not need to be specified and can generally be unrestricted, i.e., set by the bounds of the original $P=1$ action space. However, it might be beneficial to further restrict these edge parameters to get the required resolution, in cases where their optimal values vary slowly with $P$ and high circuit depths are considered.
Note that SSR does not affect the exponential scaling of the tree; in fact, after the first run, the restriction generally breaks inversion symmetry, effectively doubling the size of the decision tree, with our formulation of the problem. However, SSR makes the decision tree less complex, which is evident when $P$ is large. The regularity regions also become progressively smaller as $P$ increases, while the branching factor remains unchanged, making the discretization of the parameters finer at each step. To keep the overall computational cost under control, we will consider MCTS with a number of function evaluations growing linearly with the circuit depth, leading to an overall quadratic scaling for iterative SSR-MCTS starting from $P=1$.
III Results - 3SAT

### v3

@@ Abstract
The search for quantum algorithms to tackle classical combinatorial optimization problems has long been one of the most attractive yet challenging research topics in quantum computing. In this context, variational quantum algorithms (VQA) are a promising family of hybrid quantum-classical methods tailored to cope with the limited capability of near-term quantum hardware. However, their effectiveness is hampered by the complexity of the classical parameter optimization which is prone to getting stuck either in local minima or in flat regions of the cost-function landscape. The clever design of efficient optimization methods is therefore of fundamental importance for fully leveraging the potential of VQAs. In this work, we approach QAOA parameter optimization as a sequential decision-making problem and tackle it with an adaptation of Monte Carlo Tree Search (MCTS), a powerful artificial intelligence technique designed for efficiently exploring complex decision graphs. We show that leveraging regular parameter patterns deeply affects the decision-tree structure and allows for a flexible and noise-resilient optimization strategy suitable for near-term quantum devices. Our results shed further light on the interplay between artificial intelligence and quantum information and provide a valuable addition to the toolkit of variational quantum circuits.
@@ I Introduction
Overcoming the limitations of classical computing for problems hard to solve on classical machines, typically those requiring exponentially large resources, has been one of the main drives behind the development of quantum computation since the first pioneering works on quantum algorithms. Indeed, some celebrated examples, such as Grover’s algorithm for unstructured search [1], its generalization to more complex problems such as amplitude estimation [2] and minimum finding [3], and Shor’s algorithm for factorization [4] display significant speedup over classical methods, motivating further research in the quest for quantum advantage. However, the realization of fault-tolerant quantum computers, a prerequisite for fully exploiting this potential, remains a formidable challenge. In the interim, leveraging available technology necessitates exploring alternative strategies, such as variational quantum algorithms [5, 6, 7, 8] (VQA), to harness the computational power of current quantum devices.
@@
Despite the flexibility and popularity of VQAs, one of their main weaknesses appears in the classical optimization loop [35], which may suffer from hardware and measurement noise or the presence of flat regions in the cost function, known as barren plateaus [36]. Indeed, both effects either slow down the exploration of parameter space or prevent it from escaping local minima, in particular when the classical optimization algorithm relies on the gradient of the cost function. To tame these issues, gradient-free approaches have been suggested as appealing alternatives for the practical implementation of VQAs. However, even local minimization methods based on finite differences (also known as pseudo-gradient methods) suffer from the exponential scaling of resources when facing barren plateaus [37]. True gradient-free global optimizers are promising candidates to overcome such difficulties. Among these, reinforcement-learning [38] inspired approaches emerged as intriguing choices, both for the parameter optimization [29, 30, 39] and circuit design [32, 40].
@@
In this work, we investigate the potential of MCTS in identifying optimal strategies for QAOA applied to instances of the optimization versions of 3-SAT and MaxCut, both $\mathsf{NP}$-hard classical optimization problems [47]. We show that a suitably modified version of MCTS offers a compelling practical route to design robust optimal QAOA schedules, with exceptional resilience against noise in the cost function. The crucial modification to MCTS aims at exploiting the parameter concentration and regularity inherent in many optimal QAOA schedules [48, 49, 50, 27]. This version, which we denote as iterative Search-Space Restricted Monte Carlo Tree Search (SSR-MCTS), progressively builds the solutions at each depth $P$ from the previous optima, inspired by iterative methods presented in Refs. [11, 21]. On the opposite, we find that a vanilla version of MCTS spectacularly fails in navigating the QAOA energy landscape already at a small circuit depth. This happens because of the particular fractal-like structure of the leaf nodes, the endpoints of the decision tree, where good strategies for QAOA are uniformly and sparsely distributed among bad ones. Iterative SSR-MCTS drastically simplifies the leaf-node structure of the QAOA decision tree, allowing MCTS to reliably find optimal schedules for different choices of the system size, circuit depth, or problem instance. The combination of noise resilience, gradient-free optimization and good scalability in the quantum circuit volume makes our results well-suited for applications in current quantum hardware.
@@
The rest of the paper is organized as follows. In Section II we recap the general structure of MCTS algorithms (II.1), describe the MCTS approach for QAOA (II.2) and explain the SSR variant and its iterative version (II.3). In sections III and IV we present our results on instances of 3-SAT and unweighted MaxCut on 3-regular graphs, respectively. We draw our conclusion and present future outlooks in Sec. V. Further results and analysis are presented in the appendices: in Appendix A we briefly discuss a combination of MCTS and gradient descent to construct a global optimization algorithm, in Appendix B we look at an example MCTS decision tree for QAOA, in Appendix C we test the performance of various iterative SSR-MCTS variants on noisy cost landscapes, and in Appendix D we study the resource cost of our algorithm.
@@ II.1 Monte Carlo Tree Search
MCTS explores a given tree domain via random samples guided by some specified selection and playout policies, to find a path down the decision tree that maximizes a reward associated with the task at hand [41, 43]. Starting from the root node, it incrementally builds an asymmetric replica of the parts of the tree (branches) it deems most promising. The algorithm repeatedly runs a four-stage cycle, learning about the tree by growing the part stored in memory. At the same time, it keeps track and continuously updates the cumulative score $w_{i}$ of each stored node $i$ and the number of times it has been visited by the algorithm $n_{i}$. The cycles stop when the allotted computational budget is exceeded, at which point the algorithm makes a decision and suggests a strategy to follow. This usually corresponds either to the child node with the maximum average score $w_{i}/n_{i}$, or the one with the highest visit counts $n_{i}$. The larger the number of cycles, the more complete the reconstruction of the tree will be, leading to more accurate statistics of the nodes and improving the quality of the final choice. If the number of cycles is unbounded, MCTS is eventually equivalent to a brute force search and the policy converges to the optimal strategy [44].
The MCTS cycle.
@@
    In the selection stage, starting from the root, the algorithm travels the part of the tree stored in memory by selecting for each node its child that maximizes a specified selection policy. The most common choice is the UCT (Upper Confidence Bounds applied to Trees) policy [41, 51], which selects node $a^{*}$ among the children of node $i$, $A(i)$, such that
    $a^{*}=\mathrm{arg}\max\limits_{a\in A(i)}\left\{\frac{w_{a}}{n_{a}}+C\sqrt{\frac{2\ln n_{i}}{n_{a}}}\right\},$ (2)
@@ II.2 MCTS for quantum optimization
Using MCTS for QAOA parameter optimization requires formulating the task as a sequential decision problem. A straightforward way to achieve this is to build a decision tree of depth $2P$ and branching factor $b$. At each depth $i\in[1,2P]$ the algorithm must choose a value for the parameter $\theta_{i}$, where $\bm{\theta}=\{\gamma_{1},\beta_{1},...,\gamma_{P},\beta_{P}\}$, out of $b$ discrete options, $\theta_{i}\in\{\theta_{i}^{(1)},...,\theta_{i}^{(b)}\}$, and the range of the parameters depends on the spectrum and the symmetries of the Hamiltonians generating the parameterized unitary operators. As we consider Hamiltonians with an integer spectrum, the variational parameters belong to the compact interval $[0,2\pi)$. Moreover, since the cost function is real, it satisfies $F_{P}(-\bm{\gamma},-\bm{\beta})=F_{P}(\bm{\gamma},\bm{\beta})$, meaning that, if the different $\theta_{i}^{(k)}$ are evenly spaced, for every path $\bm{\theta}$ there is an equivalent one where $\theta_{i}^{(k)}\rightarrow 2\pi-\theta_{i}^{(k)}$. Hence, we can restrict the domain of the first parameter to $[0,\pi]$, effectively halving the size of the tree to a total of $\left\lceil\frac{1}{2}b^{2P}\right\rceil$ leaf nodes, without changing the discretization step.
This approach transforms a continuous optimization problem into a discrete one; however, we do not expect it to affect QAOA performances significantly. Indeed, the algorithm’s efficiency is limited by the complexity of the energy landscape and the structure of the decision tree, not by the impossibility of exactly pinpointing the target minimum because it does not fall on the discrete grid. Yet, pathological cases might exist where extremely thin minima [22, 56] and wide flat regions require more clever discretization schemes than the regular ones.
@@
As we rely on the variational principle for the optimization, the reward collected at each leaf node must be related to the energy associated with the corresponding choice of parameters $F_{P}(\bm{\gamma},\bm{\beta})$. In this work, we chose a nonlinear mapping $r(\bm{\theta})=e^{-\nu F_{P}(\bm{\theta})}$ for the rewards, with $\nu$ chosen such that the reward starts rising when energies at or below the typical ranges are considered, and the algorithm can properly gauge the goodness of the parameters in the context of the problem at hand  [^1] To ensure the theoretical bounds on the regret of the playing agent [105], the UCT selection policy with $C=\sqrt{2}$ requires the rewards to lie in [0,1], although this is not strictly necessary for MCTS to perform well, and should be studied on a case-by-case basis [53]..
In MCTS language, optimizing the QAOA parameters is a single-player game, meaning that no opposing agent is trying to reach a different goal. This has some consequences for the optimal strategy choice: in a two-player game, a meaningful strategy is based on the assumption that the opponent is trying to win as well, and high-scoring paths based on counterproductive choices of the other agent are usually not worth exploring [44]. This is not true in a single-player game, where the goal is maximizing the score considering all eligible strategies. In view of this, we implement some changes to MCTS proposed in Ref. [58]. We first modify the criterion with which MCTS decides the final move to play, by choosing the child that leads to the path of the random simulation that obtained the highest reward. Second, we expand the highest-scoring path of each turn all the way to the leaf node in between turns, thus ensuring that MCTS memorizes such good paths in their entirety. The search still follows the UCT selection policy in (2) for the selection stage, which can under or overestimate the potential of exploration of certain regions of the search domain. But with these changes, we have the certainty that the final parameters MCTS chooses will be at least as good as those in the best memorized sequence. In the following, we will denote this modified variant as MCTS-SP (single-player).
In the following sections, however, we will still consider the standard version of MCTS. First, it is interesting to see how it compares to the single-player variant and, second, we expect the single-player modifications to no longer be beneficial in some particularly relevant cases, such as when considering a noisy cost landscape (see Appendix C). In this noisy situation, the task loses its deterministic nature, and the memorized parameter sequences will likely contain inaccurate expected rewards, making the standard final choice criteria more appropriate, and rendering path memorization irrelevant.
@@ II.3 Iterative Search-Space Restricted MCTS
MCTS has been proven to work well for optimizing annealing schedules [39], although gradient-descent algorithms perform slightly better in idealized settings (noiseless and smooth cost functions), thanks in part to the continuous parameter spaces [45] [^2] We expect the real potential of MCTS as an optimizer to lie in robust handling of noisy or complex cost-landscapes where gradient-descent optimizers are prone to get stuck in local minima.. However, in our study, we found that standard MCTS is unsuited for QAOA as soon as $P>1$. The modified single-player version obtains improved results but quickly fails to learn good strategies too, even in shallow circuits, as we will discuss later on in Sec. III. Therefore, MCTS needs to be adapted further to perform well in QAOA parameter optimization, leveraging on the structure of its energy landscape.
@@
In this section, we describe how to modify the standard MCTS algorithm to successfully tackle QAOA. The core idea is not to modify the MCTS algorithm per se, but to progressively modify the space of eligible strategies. We call this method iterative search-space restriction (SSR). It consists of constraining the action space of the decision tree nodes as the circuit depth $P$ increases, according to the knowledge of previous solutions for $P-1$. Indeed, QAOA optima often display regular structures and patterns in the optimal parameters [11, 21, 30, 49], look ahead to Fig. 4(b) for an example. This property stems from the presence of optimal discretized quantum annealing schedules, which are typically robust with respect to small parameter changes in the Hamiltonian, system size, and number of time steps. These regular patterns appear as a monotonic increase (decrease) of the $\gamma_{i}$ ($\beta_{i}$) with the discrete-time index $i$, approaching a continuous function as $P\to\infty$. Hence, when increasing the number of layers, the new optimal parameters typically resemble a finer interpolation of the optimal schedule for the previous circuit depth.
@@
This method can be employed to look for the optimal QAOA angles one layer deeper when we know the optima for the previous depth, and its iterative application starting from $P=1$ allows for a systematic restriction of the action space by gradually increasing the number of parameters, without any prior knowledge of the optimal schedule. In this way, the parameters suggested by SSR-MCTS at depth $P$ are used to determine how the $P+1$ search is restricted. The different $\delta_{P}$ for each step can be adjusted to relax the restriction and help the algorithm correct its potential inaccuracies in previous circuit depths  [^3] To make the method fully adaptive, one might introduce automatic ways to control the softening constants during the restriction procedure, which within the present implementation have to be set manually.. In this iterative case, the boundaries of the first and last parameters ($i=0$ and $i=P+1$) do not need to be specified and can generally be unrestricted, i.e., set by the bounds of the original $P=1$ action space  [^4] In cases where their optimal values vary slowly with $P$ and high circuit depths are considered, the restricted regions of the first and last parameters would shrink much slower relative to the others, potentially causing them to lose the required resolution to select a value close to the optimal one. This could be easily circumvented by increasing (decreasing) the first (last) boundary values for the edge parameters to lie at most a fixed distance away from the previous optima, e.g., $\gamma^{*}_{P+1}$ would become $\gamma^{*}_{P+1}-\varepsilon(\gamma^{*}_{P+1}-\gamma^{*}_{P})$, for some $\varepsilon<1$.. This procedure is independent of the particular Hamiltonian instance $H_{C}$ one is facing: as the number of QAOA layers increases, it progressively generates a restricted action space tailored to the problem at hand, learning the structure of the parameter space from previous iterations.
As we will see in subsequent sections, SSR-MCTS makes the energy landscape seen by the MCTS agent easier to navigate, as regions of the parameter space that are uninteresting or even hindering the optimization task are gradually discarded. The complexity reduction of both the decision tree and the associated energy landscape is particularly evident when $P$ is large. Indeed, the regularity regions become smaller as $P$ increases, while the branching factor remains unchanged, making the discretization of the parameters finer at each step. Pictorially, it can be seen as zooming in the region around the target minimum with greater resolution at each iteration. For this reason, and despite this progressive simplification of the decision tree, SSR does not affect its exponential scaling with $P$; in fact, after the first run, the restriction generally breaks inversion symmetry, effectively doubling the number of leaf nodes, with our formulation of the problem.
To keep the overall computational cost under control, we will consider MCTS with a number of function evaluations growing linearly with the circuit depth, leading to an overall quadratic scaling for iterative SSR-MCTS starting from an unrestricted $P=1$ search.
III Results - 3-SAT